# Cody D. Schimming, C. J. O. Reichhardt, and C. Reichhardt

### v1

@@
[Uncaptioned image] We develop an approximate, analytical model for the velocity of defects in active nematics by combining recent results for the velocity of topological defects in nematic liquid crystals with the flow field generated from individual defects in active nematics. Importantly, our model takes into account the long-range interactions between defects that result from the flows they produce as well as the orientational coupling between defects inherent in nematics. We show that the model can analytically predict bound states between two $+1/2$ winding number defects, effective attraction between two $-1/2$ defects, and the scaling of a critical unbinding length between ±$\pm 1/2$ defects with activity. The model also gives predictions for the trajectories of defects, such as the scattering of $+1/2$ defects by $-1/2$ defects at a critical impact parameter that depends on activity. In the presence of circular confinement, the model predicts a braiding motion for three $+1/2$ defects that was recently seen in experiments.
@@ 1 Introduction
Topological defects are present in many physical systems, including solids, superfluids and superconductors, magnetic materials, complex fluids, soft and biological materials, and the cosmos 1, 2, 3, 4, 5, 6. Although the physical properties and scales of all of these systems vary widely, the progenitor of topological defects in each case is broken symmetry. Topological defects manifest as singularities in the broken symmetry order parameter, and, as such, take the form of quantized points, lines, surfaces, or hypersurfaces. The quantization of topological defects thus allows for an alternative picture of the structure and dynamics of these systems; namely, viewing topological defects as interacting particles. For some systems, such as vortices in type-II superconductors or Skyrmions in chiral magnets, the “particle model” approach has been successfully used to describe the behavior of the system 7, 8. In other systems, such as dislocations in solids and disclinations in liquid crystals, only recently have the required theoretical advances been made that permit defects to be treated as particles 9, 10, 11, 12, 13, 14, 15.
@@
Here, we focus on topological defects in two-dimensional active nematic liquid crystals. Active nematic liquid crystals, or simply “active nematics,” are fluids in which anisotropic constituents individually generate forces leading to collective macroscopic flows 16, 17. Active nematics are inherently out-of-equilibrium and exhibit chaotic flows on large scales 18, 19, 20, 21. They also exhibit spontaneous positive-negative defect pair production with self-advecting positive defects 22, 23, 10.
@@
Much of the theoretical and computational work surrounding active nematics has focused on a continuum model for the time dependence of the nematic tensor order parameter $\mathbf{Q}$ and the fluid velocity field $\mathbf{v}$ 16, 17. There has been recent interest, however, in the dynamics of the topological defects themselves for their role in colloidal assembly, microfluidic cargo transport, morphogenesis, logic operations, and optimal mixing 24, 25, 26, 27, 5, 28, 29, 6, 30, 31, 32, 33, 34, 35. The behavior of defects is typically inferred from solutions to the continuum equations for the order parameter, but the nonlinearity of the time evolution equations limits the progress that can be made in obtaining analytical results. There have been few studies that treat defects as particles due to the complicated interactions between defects involving their generated flows. In previous studies treating defects as particles, positive defects have been treated as active Brownian particles with long-ranged interactions 36. This treatment, however, does not include the coupling of defect orientations with relative positions, or the advection and shear flow interactions inherent in active nematics 37, 38.
@@
In this work, we combine recent analytical results for the flow field around a defect in an active nematic with an analytical framework to calculate the velocity of defects given the time dependence of the order parameter 11, 39, 15. Using these elements, we posit an analytical model for the velocity of defects in an active nematic with many defects. We show that the defect velocity may be decomposed into three parts: the Coulomb interaction between defects, advection from fluid flows generated by other defects, and deflection by shear flows generated by other defects. We analytically probe the interaction between two defects and show that the model is able to predict bound states between positive defects, effective attractions between negative defects, a critical unbinding length scale between positive and negative defects, and deflections of positive defects by negative defects. We study the predicted trajectories for multiple defects in a circularly confined system. For multiple positive defects, we identify several potential behaviors, such as stable configurations, periodic dynamics, ergodic trajectories, and defect braiding similar to that observed in recent experiments 34. We also compare the analytical model to a continuum model for the case of defect coarsening in a confined system. Finally, we summarize our results and discuss potential future extensions of the model and challenges that remain. Our work allows for a more systematic analysis of possible defect behaviors by providing analytical equations for their motion.
@@ 2.1 Defect velocity in an arbitrary flow field
where $C$ is a closed loop in the nematic and $\phi$ is the angle of $\mathbf{n}$ with respect to an arbitrary reference axis. Since $\mathbf{n}$ and $-\mathbf{n}$ represent the same state, it is possible for defects of charge $m=n/2$, where $n$ is an integer, to exist. The lowest energy nontrivial defects have charge $m=\pm 1/2$ 40. Here, we will only consider these half integer defects.
For systems with many defects, it is common to represent the nematic configuration using a rank-two tensor, $\mathbf{Q}$. In two dimensions, $\mathbf{Q}$ may be parameterized as $\mathbf{Q}=S\left[\mathbf{\hat{n}}\otimes\mathbf{\hat{n}}-(1/2)\mathbf{I}\right]$ where $S$ is a scalar representing the degree of local order. The tensor order parameter characterization has the benefit that topological defects are no longer singularities and are instead regularized so that $S$ varies continuously and $S\to 0$ at the defect cores. The apolar symmetry is also inherently present in $\mathbf{Q}$ due to its quadratic dependence on $\mathbf{n}$.
Recently, an exact expression for the velocity of defects in terms of derivatives of $\mathbf{Q}$ was derived in both two and three dimensions 11, 15. In two dimensions it reads
@@
To obtain an approximate analytical expression for the velocity of defects, we assume the Beris-Edwards model for the time evolution of $\mathbf{Q}$ in the presence of a fluid flow $\mathbf{u}$ 41:
@@
is a generalized tensor advection. Here $\mathbf{E}$ is the strain rate tensor, $\bm{\Omega}$ is the vorticity tensor and $\lambda$ is a dimensionless parameter that characterizes the tendency of nematogens to either tumble or align due to shear flow 42. The free energy, $F$, is chosen so that its bulk density is analytic in $\mathbf{Q}$ (e.g. Landau-de Gennes 40) and has a single constant elastic density $f_{E}=L|\nabla\mathbf{Q}|^{2}$. For the remainder of the paper we work in dimensionless units by scaling lengths by the nematic correlation length $\xi=\sqrt{kL/C}$ and times by the nematic relaxation time $\tau=\gamma/(C\xi^{2})$, where $k$ is a dimensionless parameter that determines the size of defects and relative strength of elastic forces and $C$ is a characteristic energy density scale of the free energy.
Equation (3) may be substituted into Eq. (2) to yield an equation for the defect velocity in terms of spatial derivatives of $\mathbf{Q}$ only. To approximate spatial derivatives at the location of the defect core, we use a linear core approximation for a nematic defect in the presence of other defects 12, 15. In the absence of flow, $\mathbf{u}=0$, the velocity of the $j$th defect reduces to
@@
where $\mathbf{r}_{ij}=\mathbf{r_{j}}-\mathbf{r_{i}}$ is the vector pointing from defect $i$ to defect $j$. Equation (5) is precisely the Coulomb interaction between defects in passive nematic liquid crystals 40. In what follows, we will set $k=4$ unless otherwise specified. In the presence of a flow field, $\mathbf{u}\neq 0$, Eq. (2) gives two additional contributions to the defect velocity. The first is advection by the flow:
@@
which we assume to be traceless since we will only consider incompressible flows. Then, introducing the vector $\mathbf{e}=E_{1}\mathbf{\hat{x}}+E_{2}\mathbf{\hat{y}}$, we have the last contribution to defect velocity
$\mathbf{v}_{j}^{\text{Shear}}=-\lambda\left[\left(\mathbf{e}\cdot\mathbf{\hat{p}}_{j}\right)\mathbf{\hat{x}}+2m_{j}\left(\mathbf{e}\cdot\left(\mathbf{\hat{p}}_{j}\cdot\bm{\varepsilon}\right)\right)\mathbf{\hat{y}}\right]$ (10)
where $\mathbf{e}$ is computed at $\mathbf{r}=\mathbf{r}_{j}$. We note that Eq. (10) depends on the charge of the defect and, in particular, the $y$-component of the velocity changes sign for different signed charges. For the remainder of the paper we will set $\lambda=1$.
@@ 2.2 Flow from active nematic defects
In this section we review the analytical results from Ref. 39 for the flow field around a defect in an active nematic and combine them with the results of the previous section to obtain an expression for the velocity of a defect in an active nematic. We consider solutions of the incompressible Stokes equation with active nematic stress:
$\eta\nabla^{2}\mathbf{u}-\Gamma\mathbf{u}$ $=\nabla p+\zeta\nabla\cdot\mathbf{Q}$ (11)
@@
where $\eta$ is the fluid viscosity, $\Gamma$ is the friction coefficient with the substrate, $p$ is the fluid pressure, and $\zeta$ is the strength of the active force. Dividing Eq. (11) by $\Gamma$ yields the parameters $\alpha\equiv\zeta/\Gamma$, which characterizes the magnitude of the active force and hence determines the magnitude of the flows, and $\ell_{h}\equiv\sqrt{\eta/\Gamma}$, which is the hydrodynamic screening length that characterizes the spatial extent of the flows.
We will assume that the total flow at the core of a defect is the sum of all flow fields created by defects. We justify this assumption by first noting that the director field in a system of many defects is determined by the Euler-Lagrange equation $\nabla^{2}\phi=0$. This equation is linear, so the full solution is the sum of each individual defect solution. In particular, we will assume the director structure is given by the minimal energy solution, so that there are no torques on defects due to elastic forces 43, 44. Physically, this corresponds to assuming that the nematic relaxation time is much smaller than the active force time scale. The flow field created by a defect is sourced by the term $\alpha\nabla\cdot\mathbf{Q}$ in the rescaled Eq. (11). Away from the core of other defects, this term depends only on the director, and, under our linear core assumptions, is given by the sum of all other defects. We thus approximate the total active force by the sum of active forces produced by each defect. This approximation is better when defects are well separated compared to their core size, and will begin to break down as defect cores overlap. Nevertheless, these approximations and assumptions allow us to make analytical progress in treating interactions between active nematic defects.
Fig.  2: (a) Radial parts $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ of the flow generated by active nematic defects plotted vs $r/2l_{h}$. (b) Radial parts $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$ of the shear rate generated by active nematic defects plotted vs $r/2l_{h}$.
In active nematics, $+1/2$ defects are much more motile than $-1/2$ defects. This is because they generate a nonzero flow at their core 39
@@
On the other hand, $-1/2$ defects generate zero flow at their core, $\mathbf{u}^{-}(0)=0$, though they still move due to interactions with other defects and flows. The flow field away from the core of ±$\pm 1/2$ defects is given by 39
@@
The components of $\mathbf{e}$, $E_{1}$ and $E_{2}$, may be computed from gradients of the velocity. For ±$\pm 1/2$ defects these are given by
@@
are plotted in Fig. 2(b). For large arguments, $\chi_{i}(x)\sim 1/x^{2}$, and hence the shear flow interaction will be subdominant when distances are much larger than the hydrodynamic screening length. This is the intuitive result for “dry” active nematics where $\ell_{h}\to 0$. In this case, the self-propulsion of $+1/2$ defects dominates other terms in the velocity. When $r\lessapprox\ell_{h}$, however, hydrodynamic interactions are much more important and complicate the motion of defects.

### v2

@@
[Uncaptioned image] We develop an approximate, analytical model for the velocity of defects in active nematics by combining recent results for the velocity of topological defects in nematic liquid crystals with the flow field generated from individual defects in active nematics. Importantly, our model takes into account the long-range interactions between defects that result from the flows they produce as well as the orientational coupling between defects inherent in nematics. Our work complements previous studies of active nematic defect motion by introducing a linear approximation that allows us to treat defect interactions as two-body interactions and incorporates the hydrodynamic screening length as a tuning parameter. We show that the model can analytically predict bound states between two $+1/2$ winding number defects, effective attraction between two $-1/2$ defects, and the scaling of a critical unbinding length between ±$\pm 1/2$ defects with activity. The model also gives predictions for the trajectories of defects, such as the scattering of $+1/2$ defects by $-1/2$ defects at a critical impact parameter that depends on activity. In the presence of circular confinement, the model predicts a braiding motion for three $+1/2$ defects that was recently seen in experiments, as well as stable and ergodic trajectories for four or more defects.
@@ 1 Introduction
Topological defects are present in many physical systems, including solids, superfluids and superconductors, magnetic materials, complex fluids, soft and biological materials, and the cosmos 1, 2, 3, 4, 5, 6. Although the physical properties and scales of all of these systems vary widely, the progenitor of topological defects in each case is broken symmetry. Topological defects manifest as singularities in the broken symmetry order parameter, and, as such, take the form of quantized points, lines, surfaces, or hypersurfaces. The quantization of topological defects thus allows for an alternative picture of the structure and dynamics of these systems; namely, viewing topological defects as interacting particles. For some systems, such as vortices in type-II superconductors or Skyrmions in chiral magnets, the “particle model” approach has been successfully used to describe the behavior of the system 7, 8. In other systems, such as dislocations in solids and disclinations in liquid crystals, only recently have the required theoretical advances been made that permit defects to be treated as particles by relating them to gradients of the continuum order parameter. 9, 10, 11, 12, 13, 14, 15, 16.
@@
Here, we focus on topological defects in two-dimensional active nematic liquid crystals. Active nematic liquid crystals, or simply “active nematics,” are fluids in which anisotropic constituents individually generate forces leading to collective macroscopic flows 17, 18. Active nematics are inherently out-of-equilibrium and exhibit chaotic flows on large scales 19, 20, 21, 22. They also exhibit spontaneous positive-negative defect pair production with self-advecting positive defects 23, 24, 11.
@@
Much of the theoretical and computational work surrounding active nematics has focused on a continuum model for the time dependence of the nematic tensor order parameter $\mathbf{Q}$ and the fluid velocity field $\mathbf{v}$ 25, 17, 18. There has been recent interest, however, in the dynamics of the topological defects themselves for their role in colloidal assembly, microfluidic cargo transport, morphogenesis, logic operations, and optimal mixing 26, 27, 28, 29, 5, 30, 31, 6, 32, 33, 34, 35, 36, 37. The behavior of defects is typically inferred from solutions to the continuum equations for the order parameter, but the nonlinearity of the time evolution equations limits the progress that can be made in obtaining analytical results, though recent progress has been made by using low-order expansions of the equations of motion of $\mathbf{Q}$ 38, 39.
There have been considerable efforts to understand the flow fields produced by individual defects in active nematics, both with and without substrate friction 23, 40, 41, 20, 42, 12, 43, 44. In many of these studies, defect motion is assumed to be dominated by the advection of the defect by the flows they produce. Additionally, because of the reliance on exact solutions for the fluid flows, understanding the motion of multiple interacting defects through them remains a challenge. Other recent studies of defect dynamics on curved surfaces have made the simplifying assumptions that positive defects are self-propelled, while negative defects are passive and only interact through Coulomb potentials and surface curvature coupling 45, 46. These models introduce noise (an experimental reality) and allow for independent orientations of defects, however, they do not include the coupling of defect orientations with relative positions, or the advection and shear flow interactions inherent in active nematics 47, 48. Recent work has also explored the effect of orientational noise on the unbinding of defects24 and developed a hydrodynamic description of many interacting defects 11, 12, 35. The hydrodynamic descriptions are suitable for systems with large numbers of active defects in the turbulent regime, but do not give descriptions of the trajectories of individual defects.
@@
In this work, we combine recent analytical results for the flow field around a defect in an active nematic with an analytical framework to calculate the velocity of defects given the time dependence of the order parameter 12, 43, 16. Using these elements, we posit an analytical model for the velocity of defects in an active nematic with many defects. A critical component to our model is a linear perturbation approximation for the $\mathbf{Q}$-tensor order parameter near defect cores. Using this, we show that the defect velocity results from two-body interactions with other defects, and may be decomposed into three parts: the Coulomb interaction between defects, advection from fluid flows generated by other defects, and deflection by shear flows generated by other defects. We analytically probe the interaction between two defects and show that the model is able to predict phenomena observed in continuum simulations and experiments, and predicted by other analytical models such as bound states between positive defects and deflections of positive defects by negative defects 12, 49. We also show that the model predicts an orientation dependent, effective attraction between negative defects and that the critical unbinding length scales proportional to the active length, both of which have not been shown in previous analytical models to our knowledge. We study the predicted trajectories for multiple defects in a circularly confined system and identify several potential behaviors, such as stable configurations, periodic dynamics, ergodic trajectories, and defect braiding similar to that observed in recent experiments 36. We also compare the analytical model to a continuum model for the case of defect coarsening in a confined system. Finally, we summarize our results and discuss potential future extensions of the model and challenges that remain. Our work complements the previous efforts to understand active nematic defect motion by explicitly including the hydrodynamic screening length as a tuning parameter, and allowing for the study of multiple interacting defects through a key approximation.
@@ 2.1 Defect velocity in an arbitrary flow field
where $C$ is a closed loop in the nematic and $\phi$ is the angle of $\mathbf{n}$ with respect to an arbitrary reference axis. Since $\mathbf{\hat{n}}$ and $-\mathbf{\hat{n}}$ represent the same state, it is possible for defects of charge $m=n/2$, where $n$ is an integer, to exist. The lowest energy nontrivial defects have charge $m=\pm 1/2$ 50. Here, we will only consider these half integer defects.
For systems with many defects, it is common to represent the nematic configuration using a rank-two tensor, $\mathbf{Q}$. In two dimensions, $\mathbf{Q}$ may be parameterized as $\mathbf{Q}=S\left[\mathbf{\hat{n}}\otimes\mathbf{\hat{n}}-(1/2)\mathbf{I}\right]$ where $S$ is a scalar representing the degree of local order. The tensor order parameter characterization has the benefit that topological defects are no longer singularities and are instead regularized so that $S$ varies continuously and $S\to 0$ at the defect cores. The apolar symmetry is also inherently present in $\mathbf{Q}$ due to its quadratic dependence on $\mathbf{\hat{n}}$.
Recently, an exact expression for the velocity of defects in terms of derivatives of $\mathbf{Q}$ was derived in both two and three dimensions 12, 16. The derivation, of which we give more details in the Appendix, is based on the conservation of topological charge for systems with a continuous, broken symmetry order parameter, first explored by Halperin51 and then later developed by Mazenko52, 53, 54. In two dimensions it reads
@@
To obtain an approximate analytical expression for the velocity of defects, we assume the Beris-Edwards model for the time evolution of $\mathbf{Q}$ in the presence of a fluid flow $\mathbf{u}$ 55:
@@
is a generalized tensor advection. Here $\mathbf{E}$ is the strain rate tensor, $\bm{\Omega}$ is the vorticity tensor and $\lambda$ is a dimensionless parameter that characterizes the tendency of nematogens to either tumble or align due to shear flow 56. The free energy, $F$, is chosen so that its bulk density is analytic in $\mathbf{Q}$ (e.g. Landau-de Gennes 50) and has a one-constant elastic density $f_{E}=L|\nabla\mathbf{Q}|^{2}$. For the remainder of the paper we work in dimensionless units by scaling lengths by the nematic correlation length $\xi=\sqrt{kL/C}$ and times by the nematic relaxation time $\tau=\gamma/(C\xi^{2})$, where $k$ is a dimensionless parameter that determines the size of defects and relative strength of elastic forces and $C$ is a characteristic energy density scale of the free energy.
Equation (3) may be substituted into Eq. (2) to yield an equation for the defect velocity in terms of spatial derivatives of $\mathbf{Q}$ only. To approximate spatial derivatives at the location of the defect core, we use a linear core approximation for a nematic defect in the presence of other defects 13, 16. In the absence of flow, $\mathbf{u}=0$, the velocity of the $j$th defect reduces to
@@
where $\mathbf{r}_{ij}=\mathbf{r}_{j}-\mathbf{r}_{i}$ is the vector pointing from defect $i$ to defect $j$. Equation (5) is precisely the Coulomb interaction between defects in passive nematic liquid crystals 50. In what follows, we will set $k=4$ unless otherwise specified. In the presence of a flow field, $\mathbf{u}\neq 0$, Eq. (2) gives two additional contributions to the defect velocity. The first is advection by the flow:
@@
which we assume to be traceless since we will only consider incompressible flows. Then the last contribution to defect velocity is
$\mathbf{v}_{j}^{\text{Shear}}=-\lambda\left[\left(E_{1}\cos 2\varphi_{j}+E_{2}\sin 2\varphi_{j}\right)\mathbf{\hat{x}}\right.\\ \left.+2m_{j}\left(-E_{1}\sin 2\varphi_{j}+E_{2}\cos 2\varphi_{j}\right)\mathbf{\hat{y}}\right]$ (10)
where $E_{1}$ and $E_{2}$ are computed at $\mathbf{r}=\mathbf{r}_{j}$. We note that Eq. (10) depends on the charge of the defect and, in particular, the $y$-component of the velocity changes sign for different signed charges. For the remainder of the paper we will set $\lambda=1$.
@@ 2.2 Flow from active nematic defects
In this section we review the analytical results from Ref. 43 for the flow field around a defect in an active nematic and combine them with the results of the previous section to obtain an expression for the velocity of a defect in an active nematic. We consider solutions of the incompressible Stokes equation with active nematic stress:
$\eta\nabla^{2}\mathbf{u}-\Gamma\mathbf{u}$ $=\nabla p+\tilde{\alpha}\nabla\cdot\mathbf{Q}$ (11)
@@
where $\eta$ is the fluid viscosity, $\Gamma$ is the friction coefficient with the substrate, $p$ is the fluid pressure, and $\tilde{\alpha}$ is the strength of the active force. Dividing Eq. (11) by $\Gamma$ yields the parameters $\alpha\equiv\tilde{\alpha}/\Gamma$, which characterizes the magnitude of the active force and hence determines the magnitude of the flows, and $\ell_{h}\equiv\sqrt{\eta/\Gamma}$, which is the hydrodynamic screening length that characterizes the spatial extent of the flows. We note that (as is done in Ref. 43) we have neglected the effect of elastic stresses in Eq. (11). To include these terms in our model, we would require an analytical solution to Eq. (11) that includes them, which, to our knowledge, has not been found. Nevertheless, it has been shown that the active stress dominates elastic stresses in other solutions of the Stokes equation for nematic defects 23, 40, 41. In a passive nematic, it would likely be much more important to include these terms, as the elastic backflow has been shown to alter the interaction of defects when active stresses are not present 57.
We will assume that the total flow at the core of a defect is the sum of all flow fields created by defects. We justify this assumption by first noting that the director field in a system of many defects is determined by the Euler-Lagrange equation $\nabla^{2}\phi=0$. This equation is linear, so the full solution is the sum of each individual defect solution. In particular, we will assume the director structure is given by the minimal energy solution, so that there are no torques on defects due to elastic forces 58, 9. Physically, this corresponds to assuming that the nematic relaxation time is much smaller than the active force time scale. The flow field created by a defect is sourced by the term $\alpha\nabla\cdot\mathbf{Q}$ in the rescaled Eq. (11). Away from the core of other defects, this term depends only on the director, and, under our linear core assumptions, is given by the sum of all other defects. We thus approximate the total active force by the sum of active forces produced by each defect. This approximation is better when defects are well separated compared to their core size, and will begin to break down as defect cores overlap. Nevertheless, these approximations and assumptions allow us to make analytical progress in treating interactions between multiple active nematic defects, since they allow us to treat the velocities as arising from two-body interactions.
Fig.  2: (a) Radial parts $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ of the flow generated by active nematic defects plotted vs $r/2\ell_{h}$. Inset: Log-log plot of $\beta_{1}$ vs $\pi/4+r/2\ell_{h}$. (b) Radial parts $\chi_{1}$, $\chi_{2}$, and $\chi_{3}$ of the shear rate generated by active nematic defects plotted vs $r/2\ell_{h}$.
In active nematics, $+1/2$ defects are much more motile than $-1/2$ defects. This is because they generate a nonzero flow at their core 43
@@
On the other hand, $-1/2$ defects generate zero flow at their core, $\mathbf{u}^{-}(0)=0$, though they still move due to interactions with other defects and flows. The flow field away from the core of ±$\pm 1/2$ defects is given by 43
@@
$E_{1}$ and $E_{2}$ may be computed from gradients of the velocity. For ±$\pm 1/2$ defects these are given by
@@
are plotted in Fig. 2(b). For large arguments, $\chi_{i}(x)\sim 1/x^{2}$, and hence the shear flow interaction will be subdominant when distances are much larger than the hydrodynamic screening length. This is the intuitive result for “dry” active nematics where $\ell_{h}\to 0$. In this case, the self-propulsion of $+1/2$ defects dominates other terms in the velocity.
We note that the limit $\ell_{h}\to\infty$ is not a sensible limit to take in the model. This is because, in this limit, the substrate friction goes to zero and solutions for the two-dimensional Stokes equation in an infinite domain diverge. This does not contradict experiments because substrate friction always plays a role, or in continuum simulations that set $\Gamma=0$ since the finite size of domains always acts as an effective screening length on the system. A more sensible limit to examine in our model is how the flow solutions behave in the limit $r<<\ell_{h}$. While the form of $\beta_{i}$ and $\chi_{i}$ is very complicated in this limit, we use the first two terms of $\beta_{1}$ (given in the Appendix) to estimate that $\beta_{1}(x)\sim(\pi^{2}/4)/(\pi/4+x)$ for small $x$. Indeed, for small $r/2\ell_{h}$ we find this to be a satisfactory approximation, as we show in the inset of Fig. 2(a).